OpenStudy (samigupta8):

The number of solutions of the equation, Arcsinx=2arctanx (in principal values) is

OpenStudy (samigupta8):

My ans is 1 bt it"z not right...

OpenStudy (samigupta8):

@michele_laino

OpenStudy (samigupta8):

@astrophysics

OpenStudy (samigupta8):

@directrix @hero

OpenStudy (priyar):

is the answer 3 ?

OpenStudy (priyar):

x=n(pi) where n=0,1,2 since principal values. i think u must have considered only zero..! right?

OpenStudy (samigupta8):

How do u know the value of arc tan π N even x can be in between [-1,1] The domain of sin inverse x is this only .. U took x=pi which is never possible

OpenStudy (samigupta8):

@mayankdevnani

OpenStudy (michele_laino):

here is my reasoning: we can write this: \[\Large \sin \theta = x,\quad \tan \left( {\frac{\theta }{2}} \right) = x\] furthermore: \[\Large \cos \theta = \sqrt {1 - {x^2}} \]

OpenStudy (samigupta8):

Sir what I did was I converted both lhs and rhs in the form of arctanx

OpenStudy (samigupta8):

Like rhs would be arctan 2x/1-x^2 And lhs would be arctanx/√1-x^2

OpenStudy (michele_laino):

yes! correct, since by substitution, we get this equation: \[\Large \tan \theta = \frac{{2\tan \left( {\theta /2} \right)}}{{1 - {{\left\{ {\tan \left( {\theta 2} \right)} \right\}}^2}}} = \frac{{2x}}{{1 - {x^2}}}\]

OpenStudy (michele_laino):

and therefore: \[\Large \begin{gathered} \frac{x}{{\sqrt {1 - {x^2}} }} = \tan \theta = \frac{{2\tan \left( {\theta /2} \right)}}{{1 - {{\left\{ {\tan \left( {\theta /2} \right)} \right\}}^2}}} = \frac{{2x}}{{1 - {x^2}}} \hfill \\ \hfill \\ \frac{x}{{\sqrt {1 - {x^2}} }} = \frac{{2x}}{{1 - {x^2}}} \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

Yep exactly....:)

OpenStudy (samigupta8):

When I solved this I got only 1 solution which is x=0

OpenStudy (michele_laino):

please wait, I'm checking such solution

OpenStudy (michele_laino):

we have these steps: \[\Large \begin{gathered} \frac{x}{{\sqrt {1 - {x^2}} }} = \frac{{2x}}{{1 - {x^2}}} \hfill \\ \hfill \\ \frac{x}{{\sqrt {1 - {x^2}} }} - \frac{{2x}}{{1 - {x^2}}} = 0 \hfill \\ \hfill \\ \frac{x}{{\sqrt {1 - {x^2}} }}\left( {1 - \frac{2}{{\sqrt {1 - {x^2}} }}} \right) = 0 \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

Yep right

OpenStudy (samigupta8):

Now 1-x^2=4 will not give us solution

OpenStudy (samigupta8):

Complex roots

OpenStudy (michele_laino):

the first equation, is: \[\Large \frac{x}{{\sqrt {1 - {x^2}} }} = 0\] whose solution, is \(x=0\), which is acceptable, since at x=0, the radical exists

OpenStudy (samigupta8):

Yep n we cannot find real roots from second equation

OpenStudy (samigupta8):

So and shud be 1 only...isn't it?

OpenStudy (samigupta8):

Ans*

OpenStudy (michele_laino):

that's right! second equation is an impossible equation in the set of the real numbers

OpenStudy (samigupta8):

Sir BT ans key says 3

OpenStudy (michele_laino):

the equation \(\sin \theta = x\), has two principal solutions if \(x\) is fixed

OpenStudy (samigupta8):

Sir principal value for arc sin x is [-π/2,π/2]

OpenStudy (michele_laino):